Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A mountain resort will hold its annual one-day snowboarding competition if it snows on either Saturday or Sunday, but if

Expert replies
by VJesus12 » Wed Jun 24, 2020 3:09 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A mountain resort will hold its annual one-day snowboarding competition if it snows on either Saturday or Sunday, but if it does not snow at all it will not hold the event. If there is a 70% chance that it snows on any given day at the resort, what is the probability that the event will be held?

A. 70%
B. 80%
C. 81%
D. 90%
E. 91%

[spoiler]OA=E[/spoiler]

Source: Veritas Prep
Join the discussion
Source: — Problem Solving |

VJesus12 wrote: ↑
Wed Jun 24, 2020 3:09 am
A mountain resort will hold its annual one-day snowboarding competition if it snows on either Saturday or Sunday, but if it does not snow at all it will not hold the event. If there is a 70% chance that it snows on any given day at the resort, what is the probability that the event will be held?

A. 70%
B. 80%
C. 81%
D. 90%
E. 91%

[spoiler]OA=E[/spoiler]

Source: Veritas Prep
GIVEN: P(snow) = 0.7, so P(no snow) = 0.3

Let's use the complement to solve this.
That is, P(event is held) = 1 - P(event is NOT held)

P(event is NOT held) = P(no snow on Saturday AND no snow on Sunday)
= P(no snow on Saturday) x P(no snow on Sunday)
= 0.3 x 0.3
= 0.09

So, P(event is held) = 1 - 0.09
= 0.91

Answer: E

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

VJesus12 wrote: ↑
Wed Jun 24, 2020 3:09 am
A mountain resort will hold its annual one-day snowboarding competition if it snows on either Saturday or Sunday, but if it does not snow at all it will not hold the event. If there is a 70% chance that it snows on any given day at the resort, what is the probability that the event will be held?

A. 70%
B. 80%
C. 81%
D. 90%
E. 91%

[spoiler]OA=E[/spoiler]

Solution:

We see that they will NOT hold the event only if it does not snow on Saturday and it does not snow on Sunday. Since the probability is 0.7 that it will snow on any given day, we see that the probability that it will NOT snow on any given day is (1 - 0.7) = 0.3. We can use the formula:

P(hold the event) = 1 - P(not hold the event)

P(not hold the event) = 0.3 x 0.3 = 0.09.

So P(hold the event) = 1 - 0.09 = 0.91 = 91%.

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion